# Violation of energy conditions and the gravitational radius of the proton

### v2

@@ Abstract
Using recent lattice QCD results for gravitational form factors, we show that the Wigner-Weyl transform of the QCD energy-momentum tensor is of exotic type IV (in the Hawking-Ellis classification) in the interior of the proton, implying that in this region: (i) it has no causal eigenvectors, (ii) it violates all (linear, point-wise) energy conditions, (iii) there are local Lorentz frames in which its energy density is zero. This occurs because of the nonzero $T^{0i}$ flux associated with the angular momentum form factor $J(t)$. The energy-momentum tensor turns into a type I tensor, characteristic of “ordinary matter”, at radial distances of $1-2$ proton Compton wavelengths. This transition point defines a “gravitational radius” for the proton, whose precise determination requires accurate knowledge of QCD form factors for momentum transfer beyond $-t=2$ GeV2. Imposing the validity of the averaged null energy condition (ANEC) of gravity, we derive a new model-independent inequality for the GFFs of the proton in QCD.
@@
Introduction: Gravitational form factors (GFFs) provide valuable information about the internal structure of hadrons, such as their spatial distributions of mass [1, 2], angular momentum [3], pressure, and shear forces [4, 5, 6, 7, 8, 9, 10] – for a review, see Ref. [11]. Recent measurements of such GFFs via deeply virtual Compton scattering have generated a lot of theoretical and phenomenological activity. To date, the GFFs of the proton have also been determined up to momentum transfer ${-t}=2$ GeV2 from lattice QCD [12]. The GFFs parameterize the graviton-proton vertex [13, 14] and allow the construction of the energy-momentum tensor (EMT) $T_{\mu\nu}$ of the proton [4, 5] as the Wigner transform of the QCD energy-momentum operator; the Wigner transform provides a map from Hilbert space operators to quasi-probability distributions in phase space.
@@
In semiclassical gravity, this proton EMT provides a source for Einstein’s equations of general relativity (GR) for the metric of spacetime. Alternatively, instead of reading Einstein’s equations as matter determining gravity ($g_{\mu\nu}$), one can read the equation in reverse and determine how the spacetime with the desired properties determines the EMT needed to produce it. This is the point of the so-called energy conditions (ECs), used to constrain the set of possible EMTs that can source Einstein’s equations [15]. Examples are the null energy condition (NEC), the weak energy condition (WEC), the dominant energy condition (DEC), and the strong energy condition (SEC) [15]. Though not fundamental [16], point-wise ECs have been instrumental in general relativity over the years to rigorously establish, e.g., incompleteness (“singularity”) theorems and the area law theorem for black holes [17]. Violations of ECs are widely known and, often, very consequential. SEC is a sufficient condition to ensure the focusing of timelike or null geodesics so that gravity is always attractive in GR [15], but this condition is violated during inflation and the current accelerated expansion of the universe. Other examples where point-wise ECs are violated, usually because of quantum effects, can be found in [18, 19, 20, 21]. Nevertheless, less stringent conditions obtained by imposing that ECs should only hold on average, such as the averaged null energy condition (ANEC), have played an important role in general relativity [22, 18, 23, 24, 25, 26, 27, 28] and are known to hold in unitary, Lorentz-invariant, interacting quantum field theories in Minkowski spacetime [29, 30].
[…]
@@
Gravitational form factors and the Wigner transform of the QCD energy-momentum operator: The vertex of a $2^{++}$ glueball or graviton and a proton can be parameterized in terms of three QCD form factors [13, 14, 4, 5] as
$p^{\prime},s^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p,s\right>=\bar{u}_{s^{\prime}}(p^{\prime})\left[A(t)\frac{P_{\mu}P_{\nu}}{m}+J(t)\frac{i(P_{\mu}\sigma_{\nu\rho}+P_{\nu}\sigma_{\mu\rho})\Delta^{\rho}}{2m}+D(t)\frac{\Delta_{\mu}\Delta_{\nu}-\eta_{\mu\nu}\Delta^{2}}{4m}\right]u_{s}(p)~{}.$ (1)
@@
The $A$ and $J$ FFs satisfy $A(0)=1$ so that the integral over all space of the static EMT $T_{00}$ gives the mass of the proton, Eq. (4) below. Furthermore, $J(0)=1/2$ is the spin of the proton. The $D$ FF is unconstrained except to satisfy $tD(t)\to 0$ as $t\to 0$. For asymptotic momentum transfer, $-t\to\infty$, the form factors scale as follows [32]:
@@
Note that the last expression implies a cancellation of the leading $\sim 1/t^{2}$ behavior between $A(t)$ and $2J(t)$. Besides the asymptotic scaling laws, our current theoretical knowledge of the FFs includes their determination from lattice QCD up to momentum transfer ${-t}=2$ GeV2 [12], as already mentioned. Furthermore, they have also been computed in holographic models for the proton [33, 34, 35, 36]. For the electromagnetic FFs of the pion and kaon the transition to the asymptotic scaling of perturbative QCD occurs at momentum transfer of order several GeV [37].
From the above matrix element, one can construct an energy-momentum tensor on a fixed (but arbitrary) time surface [4, 5, 9]:
$T_{\mu\nu}(\vec{x}\,)=\int\frac{\mathrm{d}^{3}\Delta}{2E\,(2\pi)^{3}}\,e^{-i\vec{\Delta}\cdot\vec{x}}\,\left<\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{\mu\nu}(0)\,\right|-\frac{\vec{\Delta}}{2},s\right>~{},$ (3)
@@
Coupling to classical Einstein gravity and energy conditions: In the context of semiclassical gravity, the energy-momentum tensor in (3) appears as a source in Einstein’s equations. The point-wise energy conditions we will be considering are [17]:
  • Null energy condition (NEC): $T_{\mu\nu}\ell^{\mu}\ell^{\nu}\geq 0$ for all null vectors $\ell^{\mu}$.
  • Weak energy condition (WEC): $T_{\mu\nu}t^{\mu}t^{\nu}\geq 0$ for all timelike vectors $t^{\mu}$. The physical meaning here is that the energy density measured by any observer is non-negative. By continuity, this will also be true for any null vector, so WEC implies NEC.
  • Dominant energy condition: (DEC): for all future-directed timelike $t^{\mu}$, the vector $T^{\mu}_{\nu}t^{\nu}$ should be a future-directed non-spacelike vector. The physical meaning of this condition is that for any observer, the local energy density appears non-negative and the local energy flow vector is non-spacelike. One can prove that DEC implies that if $T_{\mu\nu}$ vanishes in some set region of spacetime, then it also vanishes on the future Cauchy development of that region [15]. This is a necessary, but not sufficient, condition to ensure causality. Also, one can show that DEC implies WEC.
  • Strong energy condition (SEC): for all unit timelike $t^{\mu}$, $\left(T_{\mu\nu}-\frac{T}{2}\eta_{\mu\nu}\right)t^{\mu}t^{\nu}\geq 0$, where $T=\eta_{\mu\nu}T^{\mu\nu}$.
@@
In the second line, $\vec{\sigma}_{ss}=\xi^{\dagger}_{s}\vec{\sigma}\xi_{s}$, with $\xi_{s}$ the Pauli spinor of the proton. For proton spin $+\frac{1}{2}$ along the $z$-axis we have $\xi_{s}=(1,0)^{T}$ and $\vec{\sigma}_{ss}=(0,0,1)=\hat{z}$. Also, our notation is $\vec{x}=r\,\hat{r}$.
The discriminant $\Gamma$ determines the type of energy-momentum tensor according to the Hawking-Ellis classification [15]. In regions where $\Gamma>0$, the energy-momentum tensor is of type I [15, 38], one of the eigenvectors is timelike and the others are spacelike, and all eigenvalues are real. Physically, this means that the energy flux $T^{0i}$ can always be made to vanish via a local Lorentz transformation – there is always a physical observer who finds no net energy flux in any direction – so that the EMT is diagonal in a properly defined orthonormal frame [15]. EMTs describing classical matter, where concepts such as pressure and shear stress have clear physical meaning, are expected to be of type I [15, 17]. If $\Gamma=0$, the tensor is of type II, two eigenvalues are degenerate corresponding to two null eigenvectors, and this case is typically associated with classical radiation [15]. In general, in the absence of protecting symmetries, one expects that perturbations may take the system away from type II [39]. Type III EMTs cannot occur in spherical symmetry [21], so this case is not relevant here. When $\Gamma<0$, the EMT is of type IV; one finds no causal eigenvectors, and two eigenvalues have a nonzero imaginary part. Type IV cases are exotic [18, 40, 20], and their properties away from the test-field limit in Einstein’s equations were recently investigated in [41, 21]. For type IV EMTs, it is impossible for a physical observer not to experience a nonzero net energy flux. Furthermore, there are always local Lorentz frames in which the energy density is zero [15]. Additionally, one can show that $T_{\mu\nu}t^{\mu}t^{\nu}$ cannot be bounded from below for all unit timelike vectors $t^{\mu}$ [18]. Furthermore, away from the test-field case, it has been shown [41] that there is no hypersurface-orthogonal timelike Killing vector in a spacetimes sourced by type II, III, or IV EMTs. In particular, this implies that spacetimes generated by type IV EMTs cannot be static. Finally, we note that type IV EMTs violate all the point-wise ECs mentioned above, since NEC is violated.
@@
Test of the energy conditions with QCD form factors: We employ the parameterizations of the form factors recently determined from state-of-the-art lattice QCD calculations [12]. These are dipole FFs such as
@@
and similarly for $D(t)$ and $A(t)-2J(t)$. The parameters $\alpha_{i},\Lambda_{i}$ are given in the supplemental material of Ref. [12]. In that paper, it is shown that the dipole forms hold at least up to momentum transfer $\Delta=\sqrt{-t}=\sqrt{2}$ GeV. However, we do know that $D(t)$ as well as $A(t)-2J(t)$ asymptotically transition into a $\sim 1/t^{3}$ fall-off. Hence, we restrict the dipole form for these two FFs to a maximum of $\Delta^{*}$; we then transition right away to the asymptotic form with a coefficient chosen so that the FFs are continuous at $\Delta^{*}$ although the first derivative will not be, of course. This is of no concern here since the energy conditions involve integrals over the FFs. We also know that $A(t)$ is supposed to fall off $\sim 1/t^{2}$ asymptotically so we assume that (11) extends to asymptotic momentum transfer in that case.
Figure 1: Plot of the discriminant $\Gamma(r)$ assuming lattice QCD FFs [12] for momentum transfer below $\Delta^{*}$, and a transition to the asymptotic FFs at $\Delta^{*}$. Here, $\vec{x}$ is chosen perpendicular to the direction of the spin axis, so $\vec{M}^{2}>0$. Radial distance is measured in units of the Compton wavelength $\bar{\lambda}=1/m$, where $m$ is the proton mass.
We first test the necessary conditions of Eq. (6). We find that $T_{00}+P_{t}>0$ for all $r$ except possibly for $r$ very close to 00 and large $\Delta^{*}$. However, choosing $\vec{x}$ in the $x-y$ plane and the spin direction along $\hat{z}$, we find that the discriminant $\Gamma(r)$ in the “interior” region is mostly negative, see Fig. 1 (negative values are found whenever $\vec{x}$ is not perfectly aligned with the spin direction). This occurs because of the proton’s $T^{0i}$ component, determined by the $J$ form factor. That is, in the proton’s interior, the EMT is not of “ordinary” type I but of type IV except in the “tail” of the Wigner distribution of the proton. The farther out (in momentum transfer) the lattice QCD dipole FFs hold (greater $\Delta^{*}$), the greater $-\Gamma(r)$ becomes in the interior of the proton.

### v3

@@ Abstract
The energy–momentum tensor (EMT) of the proton encodes fundamental information about its mass, pressure, and shear distributions. Using recent lattice QCD data for the gravitational form factors, we show that the Breit–frame Wigner EMT may be of Hawking–Ellis type IV in the proton’s core. Such EMT violates all pointwise energy conditions and lacks a causal rest frame so that the usual mechanical picture fails at short distances. We define the gravitational radius—a new hadronic observable—marking the scale where the EMT becomes ordinary (type I) and the classical interpretation is restored. We also derive from the Averaged Null Energy Condition (ANEC) non-perturbative, model-independent QFT constraints on gravitational form factors.
I Introduction
@@
Gravitational form factors (GFFs) provide valuable information about the internal structure of hadrons, such as their spatial distributions of mass [1, 2], angular momentum [3], pressure, and shear forces [4, 5, 6, 7, 8, 9, 10, 11] – for a review, see Ref. [12]. Recent measurements of such GFFs via deeply virtual Compton scattering have generated a lot of theoretical and phenomenological activity. To date, the GFFs of the proton have also been determined up to momentum transfer t=2{-t}=2 GeV2 from lattice QCD [13]. The GFFs parameterize the graviton-proton vertex [14, 15] and allow the construction of the energy-momentum tensor (EMT) $T_{\mu\nu}$ of the proton [4, 5] as the Wigner transform of the QCD energy-momentum operator; the Wigner transform provides a map from Hilbert space operators to quasi-probability distributions in phase space.
@@
In the context of semiclassical gravity, this proton EMT provides a source for Einstein’s equations of general relativity (GR) for the metric of spacetime [^1] Of course, because Newton’s constant $G\sim\ell_{\mathrm{Planck}}^{2}$, where $\ell_{\mathrm{Planck}}$ is the Planck length, modifications in spacetime sourced by a single proton are utterly negligible at distances probed at colliders. Our point here is purely formal, as we are not advocating that there are any relevant, observable gravitational effects sourced by the proton.. Alternatively, instead of reading Einstein’s equations as matter determining gravity ($g_{\mu\nu}$), one can read the equation in reverse and determine how the spacetime with the desired properties determines the EMT needed to produce it. This is the point of the so-called energy conditions (ECs), used to constrain the set of possible EMTs that can source Einstein’s equations [16]. Examples are the null energy condition (NEC), the weak energy condition (WEC), the dominant energy condition (DEC), and the strong energy condition (SEC) [16]. Though not fundamental [17], point-wise ECs have been instrumental in general relativity over the years to rigorously establish, e.g., incompleteness (“singularity”) theorems and the area law theorem for black holes [18]. Violations of ECs are widely known and, often, very consequential. SEC is a sufficient condition to ensure the focusing of timelike or null geodesics so that gravity is always attractive in GR [16], but this condition is violated during inflation and the current accelerated expansion of the universe. Other examples where point-wise ECs are violated, usually because of quantum effects, can be found in [19, 20, 21, 22]. Nevertheless, less stringent conditions obtained by imposing that ECs should only hold on average, such as the averaged null energy condition (ANEC), have played an important role in general relativity [23, 19, 24, 25, 26, 27, 28, 29] and are known to hold in unitary, Lorentz-invariant, interacting quantum field theories in Minkowski spacetime [30, 31]. Thus, ECs provide an important tool for revealing the general physical properties of energy-momentum tensors even when gravitational effects are negligible, as is the case here.
The standard interpretation of the proton EMT as describing some medium, in the classical sense, with pressure and shear forces is only meaningful when the typical energy conditions hold [^2] Standard Boltzmann kinetic theory satisfies all energy conditions because the phase-space distribution function $f\geq 0$. Once quantum effects are considered by the replacement of the classical phase-space distribution by the Wigner quasi-distribution, which is not non-negative, energy condition violations can happen. However, non-positivity of the Wigner distribution may occur even for type I EMTs and is unrelated to the Hawking-Ellis classification [16].. In fact, the mass radius of the proton [5, 2], $\langle r^{2}\rangle_{\mathrm{mass}}=\int d^{3}x\,\vec{x}^{\,2}\,T_{00}(\vec{x}\,)/m$ (where $m$ is the proton mass), only has straightforward interpretation when the local energy density is non-negative for all timelike observers, i.e., if WEC holds. It is therefore interesting to test the energy-momentum tensor corresponding to the GFFs of the proton for potential violation of energy conditions. In this paper, we show that at large distances from the center of the proton, all point-wise ECs appear to be satisfied, so the dilute tails of the proton behave as ordinary matter. However, using as input recent lattice results for the GFFs, we find that near the center of the proton all point-wise ECs could be violated. A definitive answer will require measurements or lattice QCD computations of GFFs to a somewhat higher momentum transfer than is presently available. Finally, we show that ANEC leads to a new model-independent constraint for the proton and pion GFFs in QCD, which can be investigated on the lattice.
We remark that Jaffe has drawn attention to ambiguities in defining spatial charge densities in hadrons with a radius of order their Compton wavelength [32], akin to issues due to the identification of the Wigner transform of a local QFT operator with a classical phase space distribution. Also, Miller has discussed definitions of the electromagnetic “radius” of the proton [33]. However, the points raised by these authors apply the same to type I EMTs such as for a spin-0 pion, and are unrelated to the systematic and observer independent Hawking-Ellis classification [16] we pursue. Thus, it is our goal here to extract from the EMT information about the causal eigenvectors and the corresponding mechanical interpretation.
Finally, we would like to note that some of the implications of the point-wise (NEC, WEC, DEC, SEC) energy conditions on model GFFs have been considered previously in Ref. [9], assuming an unpolarized proton where the spin vector $\vec{\sigma}_{ss}$ in the expression for $T^{0i}$ in (5) is set to zero. However, the relevance of the Hawking-Ellis classification of the EMT has not been realized before, and Ref. [9] implicitly assumed the EMT to be of ordinary type I; we shall show below that for a proton with definite polarization the EMT may turn out to be of type IV in the core. Furthermore, as remarked above, in this work we derive novel model-independent constraints on the GFFs of spin 0 and spin 1/2 hadrons arising from ANEC, which have not been considered heretofore.
@@
II Gravitational form factors and the Wigner transform of the QCD energy-momentum operator
The vertex of a $2^{++}$ glueball or graviton and a proton can be parameterized in terms of three QCD form factors [14, 15, 4, 5] as
$\left<p^{\prime},s^{\prime}\left|\,\hat{T}_{\mu\nu}(0)\,\right|p,s\right>=\bar{u}_{s^{\prime}}(p^{\prime})\left[A(t)\frac{P_{\mu}P_{\nu}}{m}+J(t)\frac{i(P_{\mu}\sigma_{\nu\rho}+P_{\nu}\sigma_{\mu\rho})\Delta^{\rho}}{2m}+D(t)\frac{\Delta_{\mu}\Delta_{\nu}-\eta_{\mu\nu}\Delta^{2}}{4m}\right]u_{s}(p)~.$ (1)
@@ II Gravitational form factors and the Wigner transform of the QCD energy-momentum operator
The $A$ and $J$ FFs satisfy $A(0)=1$ so that the integral over all space of the static EMT $T_{00}$ gives the mass of the proton, Eq. (4) below. Furthermore, $J(0)=1/2$ is the spin of the proton. The $D$ FF is unconstrained except to satisfy $tD(t)\to 0$ as $t\to 0$. For asymptotic momentum transfer, $t\to\infty$, the form factors scale as follows [35]:
@@
Note that the last expression implies a cancellation of the leading $\sim 1/t^{2}$ behavior between $A(t)$ and $2J(t)$. Besides the asymptotic scaling laws, our current theoretical knowledge of the FFs includes their determination from lattice QCD up to momentum transfer t=2{-t}=2 GeV2 [13], as already mentioned. Furthermore, they have also been computed in holographic models for the proton [36, 37, 38, 39]. For the electromagnetic FFs of the pion and kaon the transition to the asymptotic scaling of perturbative QCD occurs at momentum transfer of order several GeV [40].
From the above matrix element, one can construct an energy-momentum tensor on a fixed (but arbitrary) time surface $x^{0}$ [4, 5, 9]:
$T_{\mu\nu}(\vec{x}\,)\,{=\frac{\left<x^{0},\vec{x}\,\left|\,\hat{T}_{\mu\nu}(0)\,\right|\,x^{0},\vec{x}\right>}{\left<x^{0},\vec{x}\,|\,x^{0},\vec{x}\right>}}=\int\frac{\mathrm{d}^{3}\Delta}{2E\,(2\pi)^{3}}\,e^{-i\vec{\Delta}\cdot\vec{x}}\,\left<\frac{\vec{\Delta}}{2},s\left|\,\hat{T}_{\mu\nu}(0)\,\right|-\frac{\vec{\Delta}}{2},s\right$>~, (3)
@@
III Energy conditions
The point-wise energy conditions we will be considering are [18]:
  • Null energy condition: $T_{\mu\nu}\ell^{\mu}\ell^{\nu}\geq 0$ for all null vectors $\ell^{\mu}$.
  • Weak energy condition: $T_{\mu\nu}t^{\mu}t^{\nu}\geq 0$ for all timelike vectors $t^{\mu}$. The physical meaning here is that the energy density measured by any observer is non-negative. By continuity, this will also be true for any null vector, so WEC implies NEC.
  • Dominant energy condition: for all future-directed timelike $t^{\mu}$, the vector $T^{\mu}_{\nu}t^{\nu}$ should be a future-directed non-spacelike vector. The physical meaning of this condition is that for any observer, the local energy density appears non-negative and the local energy flow vector is non-spacelike. One can prove that DEC implies that if $T_{\mu\nu}$ vanishes in some set region of spacetime, then it also vanishes on the future Cauchy development of that region [16]. This is a necessary, but not sufficient, condition to ensure causality. Also, one can show that DEC implies WEC.
  • Strong energy condition: for all unit timelike $t^{\mu}$, $\left(T_{\mu\nu}-\frac{T}{2}\eta_{\mu\nu}\right)t^{\mu}t^{\nu}\geq 0$, where $T=\eta_{\mu\nu}T^{\mu\nu}$.
@@ III Energy conditions
In the second line, $\vec{\sigma}_{ss}=\xi^{\dagger}_{s}\vec{\sigma}\xi_{s}$, with $\xi_{s}$ the Pauli spinor of the proton. For proton spin $\frac{1}{2}$ along the $z$-axis we have $\xi_{s}=(1,0)^{T}$ and $\vec{\sigma}_{ss}=(0,0,1)=\hat{z}$. Also, our notation is $\vec{x}=r\,\hat{r}$.
The discriminant $\Gamma$ determines the type of energy-momentum tensor according to the Hawking-Ellis classification [16]. In regions where $\Gamma>0$, the energy-momentum tensor is of type I [16, 41], one of the eigenvectors is timelike and the others are spacelike, and all eigenvalues are real. Physically, this means that the energy flux $T^{0i}$ can always be made to vanish via a local Lorentz transformation – there is always a physical observer who finds no net energy flux in any direction – so that the EMT is diagonal in a properly defined orthonormal frame [16]. EMTs describing classical matter, where concepts such as pressure and shear stress have clear physical meaning, are expected to be of type I [16, 18]. If $\Gamma=0$, the tensor is of type II, two eigenvalues are degenerate corresponding to two null eigenvectors, and this case is typically associated with classical radiation [16]. In general, in the absence of protecting symmetries, one expects that perturbations may take the system away from type II [42]. Type III EMTs cannot occur in spherical symmetry [22], so this case is not relevant here. When $\Gamma<0$, the EMT is of type IV; one finds no causal eigenvectors, and two eigenvalues have a nonzero imaginary part. This can occur because in Lorentzian manifolds (such as Minkowski spacetime) the inner product between vectors can have any sign, which implies that the tensor used to determine the eigenvalues, $T_{\mu}^{\nu}$, does not need to have a real spectrum or be diagonalizable; complex conjugate pairs and Jordan blocks can appear [18]. Type IV cases are rather uncommon [19, 43, 21], and their properties away from the test-field limit in Einstein’s equations were recently investigated in [44, 22]. For type IV EMTs, it is impossible for a physical observer not to experience a nonzero net energy flux. Furthermore, there are always local Lorentz frames in which the energy density is zero [16]. Additionally, one can show that $T_{\mu\nu}t^{\mu}t^{\nu}$ cannot be bounded from below for all unit timelike vectors $t^{\mu}$ [19]. Hence, type IV EMTs violate all the point-wise ECs mentioned above, since NEC is violated. This illustrates the importance of checking ECs, even in flat spacetime.
For completeness, we note that, away from the test-field case, it has been shown [44] that there is no hypersurface-orthogonal timelike Killing vector in a spacetimes sourced by type II, III, or IV EMTs. In particular, this implies that spacetimes generated by type IV EMTs cannot be static.
@@
IV Test of the energy conditions with QCD form factors
We employ the parameterizations of the form factors recently determined from state-of-the-art lattice QCD calculations [13]. These are dipole FFs such as
@@ IV Test of the energy conditions with QCD form factors
and similarly for $D(t)$ and A(t)2J(t)A(t)-2J(t). The parameters $\alpha_{i},\Lambda_{i}$ are given in the supplemental material of Ref. [13]. In that paper, it is shown that the dipole forms hold at least up to momentum transfer $\Delta=\sqrt{-t}=\sqrt{2}$ GeV. However, we do know that $D(t)$ as well as A(t)2J(t)A(t)-2J(t) asymptotically transition into a $\sim 1/t^{3}$ fall-off. Hence, we restrict the dipole form for these two FFs to a maximum of $\Delta^{*}$; we then transition right away to the asymptotic form with a coefficient chosen so that the FFs are continuous at $\Delta^{*}$ although the first derivative will not be, of course. This is of no concern here since the energy conditions involve integrals over the FFs. We also know that $A(t)$ is supposed to fall off $\sim 1/t^{2}$ asymptotically so we assume that (11) extends to asymptotic momentum transfer in that case.
Figure 1: Left: Plot of the discriminant $\Gamma(r)$ assuming lattice QCD $F$s [13] (with best fit parameters) for momentum transfer below $\Delta^{*}$, and a transition to the asymptotic $F$s at $\Delta^{*}$. Here, $\vec{x}$ is chosen perpendicular to the direction of the spin axis, so $\vec{M}^{2}>0$. Radial distance is measured in units of the Compton wavelength $\bar{\lambda}=1/m$, where $m$ is the proton mass. Right: $\Gamma(r)$ for $\Delta^{*}=2$ GeV obtained by variation of the pole masses in the individual GFFs within the uncertainties quoted in Ref. [13].
We first test the necessary conditions of Eq. (6). We find that $T_{00}+P_{t}>0$ for all $r$ except possibly for $r$ very close to 0 and large $\Delta^{*}$. However, choosing $\vec{x}$ in the xyx-y plane and the spin direction along $\hat{z}$, we find that the discriminant $\Gamma(r)$ in the “interior” region is mostly negative, see Fig. 1 (negative values are found whenever $\vec{x}$ is not perfectly aligned with the spin direction). This occurs because of the proton’s $T^{0i}$ component, determined by the $J$ form factor. That is, in the proton’s interior, the EMT is not of “ordinary” type I but of type IV except in the “tail” of the Wigner distribution of the proton. The farther out (in momentum transfer) the lattice QCD dipole FFs hold (greater $\Delta^{*}$), the greater $\Gamma(r)$ becomes in the interior of the proton.
@@
In the right panel of Fig. 1 we show various curves for $\Gamma(r)$. These were obtained by variation of the pole masses $\Lambda_{q}$, $\Lambda_{g}$ in the individual GFFs within the uncertainties quoted in Ref. [13]. For example, the curves labeled $A+$ (AA-) correspond to increasing (decreasing) both $\Lambda_{Aq}$ and $\Lambda_{Ag}$ simultaneously by the uncertainty on these parameters quoted in the supplemental material of Ref. [13]. We do not vary the $\alpha_{Aq}$, $\alpha_{Ag}$ parameters since $A(0)$, $J(0)$, and $D(0)$ correspond to conserved “charges”. The figure shows that $\Gamma(r)>0$ throughout the core of the proton could be achieved only by the most rapid allowed fall-off of $J(t)$ from $J(0)=1/2$. For $\Delta^{*}>2$ GeV, a region of negative $\Gamma$ is unavoidable even in this case.
V New constraints on GFFs from ANEC
We now go beyond the strict point-wise EC considered above and investigate also the averaged null energy condition, which states that $\int d\zeta\,T_{\mu\nu}\ell^{\mu}\ell^{\nu}\geq 0$, where the integral is taken over a complete null geodesic with affine parameter $\zeta\in(-\infty,\infty)$ and tangent vector $\ell^{\mu}=dx^{\mu}/d\zeta$. In Minkowski spacetime, one can assume without any loss of generality $x^{\mu}(\zeta)=\zeta\ell^{\mu}$, with $\ell^{\mu}=(1,n^{i})$, with $\vec{n}^{2}=1$. Using the full expression for the EMT (see Appendix D), ANEC requires $\int\mathrm{d}\zeta\,\ell^{\mu}\ell^{\nu}\,T_{\mu\nu}=\int\mathrm{d}\zeta\,\left(T^{00}-2n^{i}T^{0i}+n^{i}n^{j}T^{ij}\right)\geq 0$. For $\vec{P}=0$, imposing that ANEC holds in QCD gives the following constraint on the $A$ and $J$ GFFs (see the derivation in Appendix D)
$\int_{-\infty}^{0}dt\,\left[mA(t)-\frac{t}{4m}\left(A(t)-2J(t)\right)\right]\geq 0.$ (12)
It is interesting to note that in some holographic models [38] $A=2J$, so (12) holds for $\int\mathrm{d}t\,A(t)\geq 0$ provided the mass $m\geq 0$. In QCD, the integral in (12) is finite because $A-2J\sim 1/t^{3}$ asymptotically [35], but the t(A2J)-t(A-2J) term is not positive definite [13]. Nevertheless, we have confirmed that this inequality indeed holds for the proton EMT when the model GFFs used in our calculations of the point-wise ECs are employed. Thus, ANEC implies [^3] Besides the explicit proofs in Minkowski spacetime in [30, 31], it is known that violating ANEC would be generally catastrophic, as arguments exist stating that this could be used to build time machines [26, 27] and violate the (generalized) second law of thermodynamics [46]. Finally, we note that in Minkowski spacetime, all null lines are achronal so the ANEC condition discussed here is equivalent to the more stringent achronal averaged null energy condition [28], which is expected to hold even in curved spacetime. that (12) provides a new model-independent constraint that should be satisfied by proton GFFs in QCD.
We compare to ANEC for a spin-0 pion where the asymptotic $A(t)$ GFF takes monopole form [47, 48]. We obtain
$\begin{split}\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}}{4\pi\,|\vec{x}_{0\perp}|}+\int\frac{\mathrm{d}^{2}\Delta_{\perp}}{(2\pi)^{2}}\,e^{-i\vec{\Delta}_{\perp}\cdot\vec{x}_{0\perp}}\left[EA(-\Delta_{\perp}^{2})-\frac{[\Delta^{2}A(-\Delta^{2})]_{\infty}}{2\Delta_{\perp}}\right]\geq 0~.\end{split}$ (13)